Part A If you are driving 120 km/h along a straight road and you look to the side for 3.0s, how far do you travel during this inattentive perio Express your answer using two significant figures. d- 67 In Submit X Incorrect; Try Again: One attempt remaining
Answer #1

V = 120km/hr = 33.34m/s and t = 3s

d = Vt = 33.34×3 = 100m

Hence d = 100m